OpenStudy (anonymous):

Solve 8x - 2 = 3 (2 points) Answer x ≈ 2.52 x ≈ 0.26 x ≈ 2.74 x ≈ 0.40

OpenStudy (anonymous):

8x = 5 x = 5/8 x = .625 your options r wrong again

OpenStudy (anonymous):

The answer is not one of the options. :\

OpenStudy (anonymous):

It has to be!!!

OpenStudy (anonymous):

I'd say 0.40?

OpenStudy (anonymous):

When presented with an exponential equation such as 2x = 8, you solve for the variable by turning 8 into an exponential expression with a base of 2. 2x = 8 2x = 23

OpenStudy (anonymous):

logb xy = y logb x

OpenStudy (anonymous):

my answer is 100% correct...either ur question is wrong or ur teacher has typed something else

OpenStudy (anonymous):

thank you for your input:)

OpenStudy (anonymous):

It's machinely put

OpenStudy (anonymous):

hold u shud have included logs before lol

OpenStudy (anonymous):

a machine makes up these probs

OpenStudy (anonymous):

yeah they're logs

OpenStudy (anonymous):

\[8^{x-2} = 3\] Is this the problem?

OpenStudy (anonymous):

Ah, then the answer is x ≈ 2.52.

OpenStudy (anonymous):

Start out by using the fact that: \(x^ax^b = x^{a+x}\) \[8^{x-2} = 3\] \[8^x8^{-2}=3\] multiply both sides by \(8^2\) \[8^x = 3*8^2 = 192\] take the log base 8 on each side \[\log_8{8^x} = \log_8 192\] \[x = \log_8{192} \] rewrite the right side to base 10 or e so you can put it in your calculator \[x = \frac{\log 192}{\log 8} \approx 2.52\]

OpenStudy (anonymous):

thank you everyone!!!!!!! that helps sooooo incredibly much!!!!!!!!!!!!!!!!!